OpenStudy (kaley246):

How many zeros does this function have? y = 5x2 + 10x - 3 (Points : 4) 2 1 0

OpenStudy (kaley246):

@brusmack do u know the answer to this one

OpenStudy (anonymous):

2

OpenStudy (anonymous):

as it is a square

OpenStudy (kaley246):

How many zeros does this function have? y = 5x2 + 10x - 3 (Points : 4) 2 1 0 @HARSH123

OpenStudy (solomonzelman):

Our polynomial is \(\large\color{black}{ \displaystyle 5x^2 + 10x - 3 \\[1.0em] }\), and I am going to use the "discriminant text" (i.e. finding what your discriminant (\(\normalsize\color{black}{ \displaystyle \bf D }\)) is) \(\large\color{black}{ \displaystyle {\bf D}=(10)^2-(4)\cdot(5)\cdot(3) }\) \(\large\color{black}{ \displaystyle {\bf D}=100-60 }\) \(\large\color{black}{ \displaystyle {\bf D}=40 }\)

OpenStudy (solomonzelman):

now, I will tell you what a certain discriminant means for a quadratic equation.

OpenStudy (kaley246):

now how many zeros does this function have

OpenStudy (solomonzelman):

(Discriminant is D) \(\small\color{black}{ \displaystyle \bullet }\) When D is a perfect square that means the quadratic can be factored, and thus has 2 zeros that are integers. \(\small\color{black}{ \displaystyle \bullet }\) When D is a positive number (i.e. greater than 0) that means the quadratic has 2 zeros, but if D is NOT a perfect square, then the zeros are irrational numbers. \(\small\color{black}{ \displaystyle \bullet }\) When D is 0, you will have only 1 zero (and this zero will be an integer. \(\small\color{black}{ \displaystyle \bullet }\) When D is a negative number, you DO NOT have REAL zeros, RATHER, you have 2 imaginary zeros.

OpenStudy (solomonzelman):

[This all, is assuming integer coefficients and integer constant in the quadratic]- which you do have in your case.

OpenStudy (solomonzelman):

if you are stuck... ask whatever you don't get.